QUESTION BEGINNING AT #13


The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall Inc. common stock was $ 61 on December 31, 20Y2. Marshall Inc. Comparative Retained Earnings Statement For the Years Ended December 31, 20Y2 and 20Y1 20Y2 20Y1 Retained earnings, January 1 $ 1,312,050 $ 1,109,050 Net income 280,000 227,100 Total $ 1,592,050 $ 1,336,150 Dividends On preferred stock $ 9,100 $ 9,100 On common stock 15,000 15,000 Total dividends $ 24,100 $ 24,100 Retained earnings, December 31 $ 1,567,950 $ 1,312,050 Marshall Inc. Comparative Income Statement For the Years Ended December 31, 20Y2 and 20Y1 20Y2 20Y1 Sales $ 1,830,475 $ 1,686,550 Cost of goods sold 613,200 564,140 Gross profit $ 1,217,275 $ 1,122,410 Selling expenses $ 432,530 $ 515,010 Administrative expenses 368,455 302,460 Total operating expenses 800,985 817,470 Income from operations $ 416,290 $ 304,940 Other income 21,910 19,460 $ 438,200 $ 324,400 Other expense (interest) 120,000 66,400 Income before income tax $ 318,200 $ 258,000 Income tax expense 38,200 30,900 Net income $ 280,000 $ 227,100 Marshall Inc. Comparative Balance Sheet December 31, 20Y2 and 20Y1 Dec. 31, 20Y2 Dec. 31, 20Y1 Assets Current assets Cash $ 234,780 $ 282,860 Marketable securities 355,350 468,750 Accounts receivable (net) 321,200 299,300 Inventories 248,200 189,800 Prepaid expenses 44,420 56,570 Total current assets $ 1,203,950 $ 1,297,280 Long-term investments 1,039,155 246,822 Property, plant, and equipment (net) 1,800,000 1,620,000 Total assets $ 4,043,105 $ 3,164,102 Liabilities Current liabilities $ 415,155 $ 462,052 Long-term liabilities Mortgage note payable, 8 % $ 670,000 $ 0 Bonds payable, 8 % 830,000 830,000 Total long-term liabilities $ 1,500,000 $ 830,000 Total liabilities $ 1,915,155 $ 1,292,052 Stockholders' Equity Preferred $ 0.70 stock, $ 20 par $ 260,000 $ 260,000 Common stock, $ 10 par 300,000 300,000 Retained earnings 1,567,950 1,312,050 Total stockholders' equity $ 2,127,950 $ 1,872,050 Total liabilities and stockholders' equity $ 4,043,105 $ 3,164,102 Required: Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year. 1. Working capital $ 788,795 2. Current ratio 2.90 3. Quick ratio 2.2 4. Accounts receivable turnover 5.9 5. Number of days' sales in receivables 61.86 days 6. Inventory turnover 2.8 7. Number of days' sales in inventory 130.35 days 8. Ratio of fixed assets to long-term liabilities 9. Ratio of liabilities to stockholders' equity 10. Times interest earned 11. Asset turnover 12. Return on total assets % 13. Return on stockholders†equity % 14. Return on common stockholders†equity % 15. Earnings per share on common stock $ 16. Price-earnings ratio 17. Dividends per share of common stock $ 18. Dividend yield %
Answer #1

13) Average stockholders' equity=$(2127950+1872050)/2= $2000000

Return on stockholders' equity= (Net income-Preferred dividend)/Average stockholders' equity

= $(280000-9100)*100/2000000

= 13.5%

14) Average common stockholders' equity= $(300000+1567950+300000+1312050)/2= $1740000

Return on common stockholders' equity= (Net income-Preferred dividend)/Average common stockholders' equity

= $(280000-9100)*100/1740000

= 15.6%

15) Number of shares= $300000/10= 30000 shares

Weighted average shares outstanding= (30000+30000)/2= 30000 shares

Earnings per share on common stock= (Net income-Preferred dividend)/Weighted average shares outstanding

= $(280000-9100)/30000

= 9.03 per share

16) Price earning ratio= Market price per share/Earning price per share

= $61/9.03

= 6.8

17) Number of shares= $300000/10= 30000 shares

Dividend per share on common stock= Dividend on common stock/Number of shares

= $15000/30000

= $0.5 per share

18) Dividend yield ratio= Dividend per share on common stock/ Market price per share

= $0.5*100/61

= 0.8%